OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. (x - 3) /x = (x - 3) / (x - 6)

OpenStudy (anonymous):

x(x-3) = (x-3)(x-6) x^2-3x = x^2-9x+18 subtact x^2 and -9x from each side 6x=18

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Yes i got that part :) thanks rly helped!!